OpenStudy (anonymous):

(6 1) (5 4) the format is ax+by=c

OpenStudy (aravindg):

use two point form

OpenStudy (aravindg):

\[\large y-y_1=(x-x_1)\dfrac{(y_2-y_1)}{(x_2-x_1)}\] where \((x_1,y_1), (x_2,y_2) \) are the 2 points